Evaluate the function.
Given that f(x) = x^4 - 1.7x^2 - 1.4x, find f(-1.5).
A) 3.3375
B) 2.9
C) -0.8625
D) 6.7875

Answer by Theo(13342)

answer appears to be 3.3375.
you just replace x with -1.5 and solve
(-1.5)^4 - (1.7*((-1.5)^2)) - (1.4*(-1.5))
(-1.5)^4 = 5.0625
(1.7*((-1.5)^2)) = 3.825
(1.4*(-1.5)) = -2.1
your equation of
(-1.5)^4 - (1.7*((-1.5)^2)) - (1.4*(-1.5))
becomes
(5.0625) - (3.825) - (-2.1)
with the answer becoming
3.3375
that would be A.
it's important to keep track of the signs.
other than that, the problem is very simple to solve.